OpenStudy (anonymous):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128 i need help on this problem

OpenStudy (jhannybean):

First find your r.

OpenStudy (jhannybean):

\[\frac{8}{-2} =~?\]\[\frac{-32}{8}=~?\]

OpenStudy (jhannybean):

...etc.

OpenStudy (solomonzelman):

basically: \(\large\color{slate}{ r=(a_n ) \div (a_{n-1}) }\)

OpenStudy (jhannybean):

arithmetic sequence? or geometric sequence? @SolomonZelman

OpenStudy (solomonzelman):

geometric

OpenStudy (solomonzelman):

what do you mean, I am right!

OpenStudy (jhannybean):

Ok :) You had written arithmetic, haha

OpenStudy (solomonzelman):

oh lol

OpenStudy (anonymous):

so which one is right ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_n=a_1\times r^{n-1} }\)

OpenStudy (solomonzelman):

do you see the common ratio?

OpenStudy (jhannybean):

Yeah... to find a common ratio, divide the preceding term by the one before it.

OpenStudy (jhannybean):

Make sure you first have a common ratio between all your terms.

OpenStudy (anonymous):

i just dont know how to even start on this problem. like im horrible in math

OpenStudy (jhannybean):

Find your ratio first. \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\frac{8}{-2} =~?\]\[\frac{-32}{8}=~?\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (solomonzelman):

use anything for "r" : \(\large\color{slate}{ r=(8) \div (-2) }\) \(\large\color{slate}{ r=(-32) \div (8) }\) \(\large\color{slate}{ r=(128) \div (-32) }\) see how I am dividing a term, by the one before to find r? this is denoted as: \(\large\color{slate}{ r=(a_n) \div (a_{n-1}) }\)

OpenStudy (anonymous):

just wondering if anyone explained the common ratio here?

OpenStudy (solomonzelman):

common ratio, if anyone didn't know: is a number by which you multiply a term, to find the term right after

OpenStudy (jhannybean):

The common ratio is the term that you find that is the same for each pair in the sequence.

OpenStudy (jhannybean):

That is why we are asking you to find the common rati between each pair of terms.

OpenStudy (anonymous):

@Jhannybean I see you did

OpenStudy (solomonzelman):

Elizabeth, lost?

OpenStudy (anonymous):

Perhaps as Einstein said "make it simple, but no simpler"

OpenStudy (anonymous):

first need to find the common ratio

OpenStudy (solomonzelman):

Elizabeth, do you want me to explain everything very thoroughly?

OpenStudy (anonymous):

@SolomonZelman is the ideal person to explain

OpenStudy (solomonzelman):

that is debatable... I wouldn't necessarily say that, whether I hold like that or not. Although I will certainly attempt my best!

OpenStudy (solomonzelman):

Elizabeth, say something, please

OpenStudy (jhannybean):

So @elizabeth710 ? Are you still here?

OpenStudy (anonymous):

yes im here

OpenStudy (solomonzelman):

okay, will start from this: Do you understand what the notations I will write in blue mean? \(\large\color{blue}{ a_1 }\) , \(\large\color{blue}{ a_2 }\) , \(\large\color{blue}{ a_3 }\) \(\large\color{blue}{ r }\) , \(\large\color{blue}{ a_n }\)

OpenStudy (anonymous):

sadly no :/

OpenStudy (jhannybean):

What do you mean by your second line?

OpenStudy (solomonzelman):

got disconnected-:(

OpenStudy (solomonzelman):

Okay, so say you have the following sequence: \(\large\color{brown}{ 1,~3,~9,~27~... }\)

OpenStudy (solomonzelman):

So \(\large\color{brown}{ 1 }\) is the first term, denoted as \(\large\color{brown}{ a_1 }\) . \(\large\color{brown}{ 3 }\) is the 2nd term, denoted as \(\large\color{brown}{ a_2 }\) . \(\large\color{brown}{ 9 }\) is the 3rd term, denoted as \(\large\color{brown}{ a_3 }\) . \(\large\color{brown}{ 27 }\) is the 4th term, denoted as \(\large\color{brown}{ a_4 }\) .

OpenStudy (solomonzelman):

and on... see ? So, can you explain to me, what would \(\large\color{brown}{" a_5" }\) mean?

OpenStudy (solomonzelman):

Elizabeth, if you are not getting what I am saying, please say so...

OpenStudy (anonymous):

okay, between each term of a geometric sequence is what we call a "common ratio"

OpenStudy (anonymous):

the reason we call it a common ratio is becuase if you divide the next term, by the previous term, you will have a ratio which is "common"

OpenStudy (anonymous):

i.e divide 8 by -2 = -4, then divide -32 by 8 = -4

OpenStudy (anonymous):

This is what @SolomonZelman is trying to explain to you

OpenStudy (anonymous):

i got a_n=-2×-4